# Roles of SN and AGN Feedback in Shaping the Baryonic Content in a Wide Range of Dark Matter Halo Mass

### v2

@@ 2.2 NewCluster
[…]
In the zoom-in region, the mass resolution is $m_{\rm{DM}}=1.6\times 10^{6}\,M_{\odot}$ for dark matter and $m_{*}=2\times 10^{4}\,M_{\odot}$ for stellar particles. The spatial resolution ranges from 53 to 107 parsecs, depending on the expansion factor. The adopted cosmology follows WMAP-7 parameters: $H_{0}=70.3\,\rm{km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{m}=0.272$, $\Omega_{\Lambda}=0.728$, $\sigma_{8}=0.810$, and $n_{s}=0.967$ (Komatsu et al., 2011).
Radiative processes are modeled assuming collisional ionization equilibrium. Metal-line cooling allows gas to cool below $10^{4}\,\rm{K}$, and additional dust cooling is implemented. A uniform UV background is turned on at $z=10$, with self-shielding based on the gas density.
Star formation occurs in gas with number density $n_{\rm{H}}>5\,\rm{cm^{-3}}$ using a thermo-turbulent efficiency model. The simulation tracks nine chemical elements (H, D, C, N, O, Mg, Fe, Si, S), with yields from Type II and Type Ia supernovae and stellar winds. It adopts the Chabrier (2003) initial mass function. SN II yields are based on Kobayashi et al. (2006), while SN Ia yields come from Iwamoto et al. (1999). Each SN Ia injects $10^{51}$ erg of energy. Feedback from both SN types is implemented using the mechanical model by Kimm & Cen (2014), and stellar winds are included following Leitherer et al. (1999).
AGN feedback is modeled with a dual-mode approach (Dubois et al., 2012), where black holes inject kinetic energy (jet mode) when the Eddington ratio $f_{\rm{Edd}}<0.01$, and thermal energy (quasar mode) for $f_{\rm{Edd}}>0.01$. The Eddington ratio is capped at 1.
A complete description of the simulation is provided in Han et al. (2025b).
@@ 3.1 The Baryonic Content in DM Halos
Figure 1: Baryon fraction, normalized to the universal value ($\rm{Y}_{200}$), as a function of halo mass for the AGNeject1 (left panel) and AGNeject2 (right panel) models. Solid lines in different colors represent predictions at various redshifts, and are compared with results from our NewCluster simulation (filled circles), along with observational and simulation data from the literature (symbols as indicated in the legend). AGNeject1 shows little to no evolution across redshift, whereas AGNeject2 predicts a significant decline in baryon content with time in the range $13.411.8<\log M_{\rm{halo}}<$ 13.4. At small scales, NewCluster supports the near-constant evolution suggested by AGNeject1. On cluster scales, both models agree reasonably well with the observations by Chiu et al. (2016) and Akino et al. (2022), but fall below the values reported by Angelinelli et al. (2023). The key difference between the models lies in the deep cavity at $z=0$ (blue line) in AGNeject2, highlighting a stronger late-time depletion of baryons.
We begin our analysis in Figure 1, which shows the baryon fraction—normalized to the universal value ($\rm{Y}_{200}$)—as a function of halo mass for the AGNeject1 (left panel) and AGNeject2 (right panel) models. Solid lines represent model predictions at different redshifts, while filled circles correspond to results from our NewCluster simulation. We also include observational estimates from Chiu et al. (2016) at $z\sim 0.9$, and Akino et al. (2022) at $10<z<$ 1, along with predictions from the MAGNETICUM simulation (Dolag et al., 2016) by Angelinelli et al. (2023), with symbols as indicated in the legend.
@@
[…]
In contrast, AGNeject2 exhibits a more pronounced redshift dependence over the mass range $13.411.8<\log M_{\rm{halo}}<$ 13.4, especially between $z=1$ and the present. This results in a distinct “cavity” in the baryon fraction at $z=0$ within that range. Outside this interval, the two models predict similar baryon fractions, indicating that the differences arise primarily at intermediate halo masses and late times—consistent with a more efficient, redshift-dependent removal of baryons in AGNeject2.
Compared to MAGNETICUM, both models predict lower baryon fractions—expected given the different apertures—but they are broadly consistent with observational data from Chiu et al. (2016) and Akino et al. (2022) for halos above $\log M_{\rm{halo}}\sim 13$, even accounting for the likely upward correction needed in the observed values.
Figure 2: Hot gas fraction, normalized to the universal baryon fraction, as a function of halo mass at different redshifts (separate panels), as predicted by our models AGNeject1 (blue lines) and AGNeject2 (red lines). The predictions are compared to both observational constraints and simulation results at $z=0$ and $z=1$, and to simulations at higher redshifts. Results from our NewCluster simulation are also shown at $z\geq 1$ as filled magenta circles. Both models are broadly consistent with observations at $z=0$ and show signs of convergence at $z=1$, though they generally predict lower hot gas fractions than those found in the MAGNETICUM simulation (Angelinelli et al., 2023) across all redshifts. The main difference between the two models emerges in Milky Way–like halos (comparable to the mass range discussed in Figure 1). AGNeject1 is more efficient at ejecting hot gas down to $z=1$, while AGNeject2 becomes more efficient in the last $\sim$7 Gyr. At $z=3$, NewCluster predicts slightly lower hot gas fractions compared to the models, but overall follows the same qualitative trends. This evolution explains the deep cavity observed in the baryon fraction at $z=0$ for AGNeject2, as seen in Figure 1.
@@
To investigate the gaseous component in more detail, Figure 2 displays the normalized hot gas fraction as a function of halo mass, with predictions from AGNeject1 (blue lines) and AGNeject2 (red lines) across different redshifts (each shown in separate panels). These are compared to predictions from the NewCluster simulation and a compilation of observational data from the literature: brown stars from Vikhlinin et al. (2006), Arnaud et al. (2007), Sun et al. (2009), Rasmussen & Ponman (2009), Pratt et al. (2009), Mahdavi et al. (2013), Gonzalez et al. (2013), Lovisari et al. (2015), Eckert et al. (2016), Pearson et al. (2017), Eckert et al. (2019), Mulroy et al. (2019), and Ragagnin et al. (2022); magenta diamonds from Popesso et al. (2024); green diamonds from Chiu et al. (2016); a grey shaded region from Eckert et al. (2021); a pink shaded region from Dev et al. (2024); and predictions from the MAGNETICUM simulation.
Before delving into the results, we note an important caveat. As in Figure 1, many of the observational estimates and MAGNETICUM predictions refer to gas fractions measured within $R_{500}$, rather than $R_{200}$. Unlike the case of total baryons, however, a correction can be applied to translate $\rm{Y}_{500}$ values into their $\rm{Y}_{200}$ equivalents. To perform this conversion, we used the empirical relations provided by Popesso et al. (2024), which link $\rm{Y}_{200}$ to $R_{200}$ and $\rm{Y}_{500}$ to $R_{500}$, based on fits to their data. This allows us to compare our model predictions directly to the corrected observational estimates shown in Figure 2.
@@
Figure 3: Stellar-to-halo mass relation for central galaxies at different redshifts (panels), as predicted by our models AGNeject1 (blue lines) and AGNeject2 (red lines). At $z=0$, the predictions are compared with the SHARK2 SAM (Lagos et al., 2024), the empirical relation from Moster et al. (2013), and observational estimates from Romeo et al. (2020), Hinshaw et al. (2013), Gonzalez et al. (2013), and Kravtsov et al. (2018), as indicated in the legend. At higher redshifts ($z\geq 1$), model predictions are compared with Moster et al. (2013) and our NewCluster simulation. Both AGNeject1 and AGNeject2 show good agreement with semi-analytic, empirical, and observational constraints at $z=0$. At higher redshifts, NewCluster tends to overpredict the stellar mass in low-mass halos. Accurately reproducing this relation across cosmic time is essential, as it confirms that central galaxies acquire a realistic stellar mass. This is especially crucial in low-mass halos, where the central galaxy typically accounts for the majority of stellar mass within the virial radius.
@@
At higher redshift, in particular at $z=1$, we start to see the difference between our models: AGNeject1 predicts a lower hot gas fraction than AGNeject2 in the mass range corresponding to the cavity at the present time. This is because AGNeject1 is effective from early times down to $z=0$, whereas AGNeject2 becomes highly effective only after $z\sim 1$. However, given the constancy of the hot gas fraction on cluster scales, our predictions remain consistent with the observations by Chiu et al. (2016), although they do not match those from MAGNETICUM, which show a constant trend even at higher redshift (see bottom panels). NewCluster, on the other hand, aligns well with our models at small scales up to $z=2$, but this agreement breaks down at $z=3$.
@@
Comparing our results qualitatively with earlier studies—specifically C25 and Popesso et al. (2024)—we find similar conclusions. These authors emphasized the necessity, in numerical models, of a feedback mechanism capable of ejecting hot gas from halos without significantly altering the stellar and cold gas components. This is precisely what our models achieve, as the next plot will illustrate. Nonetheless, it remains challenging to distinguish between the two models due to the scarcity of observational data at $z>1$ in the halo mass range where the cavity begins to develop—a feature entirely driven by more efficient hot gas expulsion.
@@
One final point worth noting is that the normalized hot gas fraction at $z=0$—where observational coverage is most comprehensive—rises from roughly 0.3 at $\log M_{\rm{halo}}\sim 12$ to about 0.65 in the most massive clusters. Comparing these figures with the total baryon fractions (which include cold gas and stars), we infer that, on cluster scales, approximately 92%-93% of the baryons reside in the hot phase, with this percentage declining toward lower halo masses. As previously shown in C25, on smaller scales, both the stellar and cold gas components become more significant.
@@
We conclude this section by demonstrating that our models also accurately predict the stellar content of halos. This is shown in Figure 3, which displays the stellar-to-halo mass relation for central galaxies at various redshifts (panels), with our models color-coded as in the previous figure. At $z=0$, we compare our results to the SHARK2 SAM (Lagos et al. 2024, black dashed lines), the empirical relation by Moster et al. 2013 (gray shaded region), and observational data from Romeo et al. 2020 (brown diamonds), Hinshaw et al. 2013, Gonzalez et al. 2013, and Kravtsov et al. 2018 (magenta squares). Both of our models match these datasets and the empirical relation very well, with no visible difference between them across the entire halo mass range.
At higher redshift, we compare our models with predictions from NewCluster (magenta filled circles) and the Moster et al. (2013) relation. Our models follow the empirical relation remarkably well up to $z=3$, underscoring their ability to predict the correct stellar mass of central galaxies across all halo masses and cosmic times. NewCluster, by contrast, appears to overpredict stellar masses at all redshifts by approximately an order of magnitude. The origin of this discrepancy will be investigated in detail in Seo et al., in prep.
Focusing solely on the predictions from our SAM, both models successfully reproduce the stellar-to-halo mass relation at every redshift considered. This is important because stars are one of the three components used in computing the baryon fraction, meaning that our models accurately describe at least this component. Furthermore, given that the tension between model predictions and observations in both the baryon and hot gas fractions arises primarily at small group scales—and considering that stellar mass is a good tracer of the total stellar content at those scales, where cold gas is negligible—we conclude that the discrepancies are driven entirely by how much hot gas is expelled by feedback processes.
Again, this conclusion is in line with the arguments made in C25 and Popesso et al. (2024), which can be summarized succinctly: the physical mechanism(s) responsible for ejecting hot gas outside halos must not disturb the stellar or cold gas components. Our two models achieve this in different ways and with varying strengths over time. With this established, we now move on to investigate the two feedback mechanisms capable of ejecting hot gas—SN and AGN feedback—to determine which dominates, when, and in which halo mass regime.
@@ 3.2 Roles of SN and AGN feedback
Figure 4: Efficiency of AGN feedback in ejecting gas beyond the virial radius, shown as a function of halo mass. The efficiency is defined as the total mass of gas expelled by AGN activity up to a given redshift (integrated over all progenitors in the merger tree), normalized by the halo mass at that redshift. Predictions from our two models are shown: AGNeject1 (left panel) and AGNeject2 (right panel), with solid lines in different colors representing different redshifts, as indicated in the legend. In AGNeject1, the ejection efficiency steadily increases with cosmic time, reaching a peak in the range $1312.5<\log M_{\rm{halo}}<$ 13, with the precise peak depending on redshift. In contrast, AGNeject2 shows negligible efficiency across all halo masses until $z=1$ (green line), after which the efficiency rises sharply at $\log M_{\rm{halo}}>12$, peaking around $\log M_{\rm{halo}}\sim 13$ at $z=0$. This illustrates the fundamentally different timing and strength of AGN-driven gas ejection in the two models.
@@
Figure 6: Fraction of gas ejected beyond the virial radius by SN (solid lines) and AGN (dashed lines) feedback, as a function of halo mass and at various redshifts (indicated by different colors), for the AGNeject1 (left panel) and AGNeject2 (right panel) models. The fractions shown represent the contribution of each feedback mode relative to the total ejected gas (i.e., SN plus AGN). The figure highlights the marked differences between the two models. In AGNeject1, AGN feedback dominates over SN feedback in the halo mass range $13.511.5<\log M_{\rm{halo}}<$ 13.5, depending on redshift, with the dominance centered at $\log M_{\rm{halo}}\sim 13$ at $z=0$. In contrast, AGNeject2 shows negligible AGN ejection until $z\leq 1$, after which it becomes efficient and overtakes SN feedback in the range $13.212.3<\log M_{\rm{halo}}<$ 13.2 at $z=0$. At the high-mass end ($\log M_{\rm{halo}}>14$), SN and AGN feedback contribute comparably in both models.
@@
This divergence emphasizes that AGNeject1 gradually expels hot gas over time, while AGNeject2 remains largely inactive for about 7 Gyr and then becomes suddenly effective, even surpassing the performance of AGNeject1 between $10<z<$ 1. In summary, by $z=0$, both models are most effective within the range $1412<\log M_{\rm{halo}}<$ 14, with an efficiency of roughly 5% on cluster scales.
@@
To directly compare SN and AGN feedback within the same framework, in Figure 6 we plot their respective contributions to the total gas expelled as a function of halo mass and redshift. SN feedback is represented by solid lines, and AGN feedback by dashed lines, with different colors indicating redshift. The left and right panels correspond to the AGNeject1 and AGNeject2 models, respectively. The two models display clearly distinct behaviors: AGNeject1 shows a weak dependence on redshift, whereas AGNeject2 exhibits a strong redshift dependence, as expected from its delayed onset of AGN activity.
Notably, in both models, SN feedback dominates the total gas expulsion except within the characteristic mass ranges where AGN feedback becomes particularly efficient. Specifically, AGN feedback dominates in halos with $13.512.5<\log M_{\rm{halo}}<$ 13.5 for AGNeject1, and $13.312.2<\log M_{\rm{halo}}<$ 13.3 for AGNeject2, both at $z=0$. Thus, SN feedback is the primary driver of gas ejection in low-mass and Milky Way-like halos, while AGN feedback takes over in the subsequent mass decade, with both mechanisms contributing roughly equally on large group and cluster scales.
Our findings are qualitatively consistent with those of Ayromlou et al. (2023) (but see also the analysis in Wright et al. 2024), who analyzed three sets of cosmological hydrodynamic simulations—EAGLE (Schaye et al., 2015), IllustrisTNG (Pillepich et al., 2018), and SIMBA (Davé et al., 2019). These authors showed that feedback processes play a crucial role in redistributing gas within and beyond halo boundaries, reducing the baryon fraction inside halos and transferring hot gas outside the virial radius.
@@
To conclude this section, we now turn to a detailed comparison between the baryon fractions predicted by our models and those reported by Ayromlou et al. (2023) in their analysis.

### v4

@@ 2.2 NewCluster
NewCluster simulation (S. Han et al. 2025b) is a high-resolution cosmological zoom-in run designed to study galaxy formation in a massive cluster environment. It uses the RAMSES-yOMP code (S. Han et al., 2025a), a hybrid-parallelized version of the RAMSES hydrodynamics solver (R. Teyssier, 2002), and focuses on the evolution of a region centered on a galaxy cluster with a virial mass of approximately $5\times 10^{14}\,M_{\odot}$ at redshift $z=0$. The simulation starts from a periodic box of 100 Mpc/h/h and zooms in on a region of radius $\sim 17.7$ Mpc/h/h, corresponding to $3.5\,R_{\rm{vir}}$.
In the zoom-in region, the mass resolution is $m_{\rm{DM}}=1.6\times 10^{6}\,M_{\odot}$ for dark matter and $m_{*}=2\times 10^{4}\,M_{\odot}$ for stellar particles. The spatial resolution ranges from 53 to 107 parsecs, depending on the expansion factor. The adopted cosmology follows WMAP-7 parameters: $H_{0}=70.3\,\rm{km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{m}=0.272$, $\Omega_{\Lambda}=0.728$, $\sigma_{8}=0.810$, and $n_{s}=0.967$ (E. Komatsu et al., 2011).
Radiative processes assume collisional ionization equilibrium, with metal-line and dust cooling allowing gas to cool below $10^{4},\rm{K}$. A uniform UV background is switched on at $z=10$, with density-based self-shielding.
Star formation occurs in gas above $n_{\rm H}>5,\rm{cm^{-3}}$, following a thermo-turbulent efficiency model. The simulation tracks nine elements (H, D, C, N, O, Mg, Fe, Si, S), adopting a G. Chabrier (2003) IMF. Yields come from SN II (C. Kobayashi et al., 2006), SN Ia (K. Iwamoto et al., 1999), and stellar winds (C. Leitherer et al., 1999). Each SN Ia injects $10^{51}$ erg, with feedback implemented via the mechanical model of T. Kimm & R. Cen (2014).
AGN feedback follows the dual-mode scheme of Y. Dubois et al. (2012): jet mode (kinetic) for $f_{\rm Edd}<0.01$, and quasar mode (thermal) for $f_{\rm Edd}>0.01$, capped at unity. Further details are given in S. Han et al. (2025b).
@@ 3.1 The Baryonic Content in DM Halos
Figure 1: Baryon fraction normalized to the universal value ($\rm{Y}{200}$) as a function of halo mass for the AGNeject1 (left) and AGNeject2 (right) models. Solid lines indicate different redshifts, compared with predictions from NewCluster (filled circles) and literature data (symbols in the legend). AGNeject1 shows almost no evolution with time, while AGNeject2 predicts a strong late-time decline in baryon content between $11.8<\log M{\rm{halo}}<13.4$. On small scales, NewCluster supports the near-constant trend of AGNeject1. On cluster scales, both models agree with I. Chiu et al. (2016) and D. Akino et al. (2022), but fall below M. Angelinelli et al. (2023). The most distinctive feature is the deep $z=0$ cavity in AGNeject2, signaling enhanced baryon depletion at late times.
We begin our analysis in Figure 1, which shows the baryon fraction—normalized to the universal value ($\rm{Y}_{200}$)—as a function of halo mass for the AGNeject1 (left panel) and AGNeject2 (right panel) models. Solid lines represent model predictions at different redshifts, while filled circles correspond to results from our NewCluster simulation. We also include observational estimates from I. Chiu et al. (2016) at $z\sim 0.9$, and D. Akino et al. (2022) at $0<z<1$, along with predictions from the MAGNETICUM simulation (K. Dolag et al., 2016) by M. Angelinelli et al. (2023), with symbols as indicated in the legend.
@@
Focusing first on the comparison between our models, AGNeject1 shows little redshift evolution, with only mild variation in a order of magnitude centered around $\log M_{\rm{halo}}\sim 12.7–12.8$. At lower halo masses, the predictions from NewCluster closely follow the nearly constant trend suggested by AGNeject1. At $z=0$, the baryon fraction increases from about 0.4 at $\log M_{\rm{halo}}\sim 11.2$ to roughly 0.75 for the most massive halos, reflecting a clear correlation with halo mass.
In contrast, AGNeject2 exhibits a more pronounced redshift dependence over the mass range $11.8<\log M_{\rm{halo}}<13.4$, especially between $z=1$ and the present. This results in a distinct “cavity” in the baryon fraction at $z=0$ within that range. Outside this interval, the two models predict similar baryon fractions, indicating that the differences arise primarily at intermediate halo masses and late times—consistent with a more efficient, redshift-dependent removal of baryons in AGNeject2.
Compared to MAGNETICUM, both models predict lower baryon fractions—expected given the different apertures—but they are broadly consistent with observational data from I. Chiu et al. (2016) and D. Akino et al. (2022) for halos above $\log M_{\rm{halo}}\sim 13$, even accounting for the likely upward correction needed in the observed values.
Figure 2: Hot gas fraction normalized to the universal baryon fraction as a function of halo mass at different redshifts (separate panels), predicted by AGNeject1 (blue) and AGNeject2 (red). Results are compared with observational constraints and simulations at $z=0$ and $z=1$, and with simulations only at higher redshifts. Predictions from NewCluster are also shown at $z\geq 1$ (magenta circles). Both models broadly agree with observations at $z=0$ and show hints of convergence at $z=1$, though they predict lower fractions than MAGNETICUM (M. Angelinelli et al., 2023) at all epochs. The main difference arises in Milky Way–like halos: AGNeject1 is more effective until $z=1$, while AGNeject2 becomes dominant in the last $\sim 7$ Gyr. At $z=3$, NewCluster gives slightly lower values but follows the same trends. This explains the deep baryon fraction cavity at $z=0$ in AGNeject2 (Figure 1).
@@
To investigate the gaseous component in more detail, Figure 2 displays the normalized hot gas fraction as a function of halo mass, with predictions from AGNeject1 (blue lines) and AGNeject2 (red lines) across different redshifts (each shown in separate panels). These are compared to predictions from the NewCluster simulation and a compilation of observational data from the literature: brown stars from A. Vikhlinin et al. (2006), M. Arnaud et al. (2007), M. Sun et al. (2009), J. Rasmussen & T. J. Ponman (2009), G. W. Pratt et al. (2009), A. Mahdavi et al. (2013), A. H. Gonzalez et al. (2013), L. Lovisari et al. (2015), D. Eckert et al. (2016), R. J. Pearson et al. (2017), D. Eckert et al. (2019), S. L. Mulroy et al. (2019), and A. Ragagnin et al. (2022); magenta diamonds from P. Popesso et al. (2024); green diamonds from I. Chiu et al. (2016); a grey shaded region from D. Eckert et al. (2021); a pink shaded region from A. Dev et al. (2024); and predictions from the MAGNETICUM simulation.
Before delving into the results, we note an important caveat. As in Figure 1, many of the observational estimates and MAGNETICUM predictions refer to gas fractions measured within $R_{500}$, rather than $R_{200}$. Unlike the case of total baryons, however, a correction can be applied to translate $\rm{Y}_{500}$ values into their $\rm{Y}_{200}$ equivalents. To perform this conversion, we used the empirical relations provided by P. Popesso et al. (2024), which link $\rm{Y}_{200}$ to $R_{200}$ and $\rm{Y}_{500}$ to $R_{500}$, based on fits to their data. This allows us to compare our model predictions directly to the corrected observational estimates shown in Figure 2.
@@
It is worth noting that, although not displayed in our figures, several recent observational studies suggest that Milky Way–mass halos may host relatively high baryon fractions. For example, S. Das et al. (2023), using the thermal SZ effect, reported a non-monotonic dependence of the baryon fraction on galaxy mass, with a peak around galaxies of $\log M_{*}\sim 11$. Similarly, F. Nicastro et al. (2023) and S. Mathur et al. (2023), through X-ray absorption measurements, estimated $\log M_{\rm CGM}\sim 11$ for halos of $\log M\sim 12$, corresponding to a normalized hot gas fraction of roughly 0.6. Taken together, these studies support the presence of a declining baryon fraction with increasing halo mass in the range $\log M\sim 12-13$, consistent with the cavity predicted by AGNeject2. This agreement with independent observations reinforces the need for feedback mechanisms, such as AGN-driven hot gas ejection, in order to reproduce the full baryon cycle in galaxies and halos.
@@
Figure 3: Stellar-to-halo mass relation for central galaxies at different redshifts (panels), as predicted by AGNeject1 (blue) and AGNeject2 (red). At $z=0$, predictions are compared with the SHARK2 SAM (C. d. P. Lagos et al., 2024), the empirical relation of B. P. Moster et al. (2013), and observational estimates from A. B. Romeo et al. (2020), G. Hinshaw et al. (2013), A. H. Gonzalez et al. (2013), A. V. Kravtsov et al. (2018), and P. E. Mancera Piña et al. (2025). At $z\geq 1$, results are compared with B. P. Moster et al. (2013) and the NewCluster simulation. Both models agree well with semi-analytic, empirical, and observational constraints at $z=0$. At higher redshifts, NewCluster tends to overpredict stellar masses in low-mass halos. Matching this relation over cosmic time is crucial, especially in low-mass halos where the central galaxy contains most of the stellar mass within the virial radius.
@@
At higher redshift, in particular at $z=1$, we start to see the difference between our models: AGNeject1 predicts a lower hot gas fraction than AGNeject2 in the mass range corresponding to the cavity at the present time. This is because AGNeject1 is effective from early times down to $z=0$, whereas AGNeject2 becomes highly effective only after $z\sim 1$. However, given the constancy of the hot gas fraction on cluster scales, our predictions remain consistent with the observations by I. Chiu et al. (2016), although they do not match those from MAGNETICUM, which show a constant trend even at higher redshift (see bottom panels). NewCluster, on the other hand, aligns well with our models at small scales up to $z=2$, but this agreement breaks down at $z=3$.
@@
Comparing our results qualitatively with earlier studies—specifically C25 and P. Popesso et al. (2024)—we find similar conclusions. These authors emphasized the necessity, in numerical models, of a feedback mechanism capable of ejecting hot gas from halos without significantly altering the stellar and cold gas components. This is precisely what our models achieve, as the next plot will illustrate. Nonetheless, it remains challenging to distinguish between the two models due to the scarcity of observational data at $z>1$ in the halo mass range where the cavity begins to develop—a feature entirely driven by more efficient hot gas expulsion.
@@
One final point worth noting is that the normalized hot gas fraction at z=0z=0—where observational coverage is most comprehensive—rises from roughly 0.3 at $\log M_{\rm{halo}}\sim 12$ to about 0.65 in the most massive clusters. Comparing these figures with the total baryon fractions (which include cold gas and stars), we infer that, on cluster scales, approximately 92%-93% of the baryons reside in the hot phase, with this percentage declining toward lower halo masses. As previously shown in C25, on smaller scales, both the stellar and cold gas components become more significant.
@@
We conclude this section by showing that our models also reproduce the stellar content of halos. Figure 3 presents the stellar-to-halo mass relation for central galaxies at different redshifts, with our models color-coded as before. At $z=0$, they match the SHARK2 SAM (C. d. P. Lagos et al., 2024), the empirical relation of B. P. Moster et al. (2013), and observational data from A. B. Romeo et al. (2020), G. Hinshaw et al. (2013), A. H. Gonzalez et al. (2013), A. V. Kravtsov et al. (2018), and P. E. Mancera Piña et al. (2025). Across the full mass range, both models agree well with data and show no visible differences.
At higher redshifts, our models also follow the B. P. Moster et al. (2013) relation up to $z=3$, confirming their ability to reproduce the stellar mass of central galaxies at all times. By contrast, NewCluster overpredicts stellar masses by about an order of magnitude, an issue to be explored in Seo C., et al. (in prep).
These results demonstrate that our SAM reliably captures the stellar-to-halo mass relation at all epochs. Since stars are one of the main components in the baryon fraction, this confirms that the models describe them accurately. Given that tensions between model and data for baryon and hot gas fractions arise mostly on group scales—where cold gas is negligible and stellar mass is a good tracer—the discrepancies must be driven by how much hot gas feedback expels.
This conclusion aligns with C25 and P. Popesso et al. (2024): the physical mechanisms removing hot gas must leave the stellar and cold components unaffected. Both of our models achieve this in different ways and with varying strength over time. With this established, we now turn to the two feedback channels able to expel hot gas—SN and AGN—to assess their relative roles across mass and redshift.
@@ 3.2 Roles of SN and AGN feedback
Figure 4: Efficiency of AGN feedback in ejecting gas beyond the virial radius as a function of halo mass. The efficiency is defined as the total gas mass expelled by AGN up to a given redshift (including all progenitors), normalized by the halo mass at that epoch. Predictions from our two models are shown: AGNeject1 (left) and AGNeject2 (right), with colors indicating different redshifts. In AGNeject1, efficiency gradually rises with time, peaking between $12.5<\log M_{\rm{halo}}<13$, depending on redshift, and flattening to $\sim 5\%$ on cluster scales. By contrast, AGNeject2 remains negligible until $z=1$, then increases sharply at $\log M_{\rm{halo}}>12$, reaching a peak near $\log M_{\rm{halo}}\sim 13$ at $z=0$. This highlights the contrasting timing and strength of AGN-driven gas ejection in the two models.
@@
Figure 6: Fraction of gas ejected beyond the virial radius by SN (solid lines) and AGN (dashed lines) feedback, as a function of halo mass and at various redshifts (indicated by different colors), for the AGNeject1 (left panel) and AGNeject2 (right panel) models. The fractions shown represent the contribution of each feedback mode relative to the total ejected gas (i.e., SN plus AGN). The figure highlights the marked differences between the two models. In AGNeject1, AGN feedback dominates over SN feedback in the halo mass range $11.5<\log M_{\rm{halo}}<13.5$, depending on redshift, with the dominance centered at $\log M_{\rm{halo}}\sim 13$ at $z=0$. In contrast, AGNeject2 shows negligible AGN ejection until $z\leq 1$, after which it becomes efficient and overtakes SN feedback in the range $12.3<\log M_{\rm{halo}}<13.2$ at $z=0$. At the high-mass end ($\log M_{\rm{halo}}>14$), SN and AGN feedback contribute comparably in both models.
@@
This divergence emphasizes that AGNeject1 gradually expels hot gas over time, while AGNeject2 remains largely inactive for about 7 Gyr and then becomes suddenly effective, even surpassing the performance of AGNeject1 between $0<z<1$. In summary, by $z=0$, both models are most effective within the range $12<\log M_{\rm{halo}}<14$, with an efficiency of roughly 5% on cluster scales.
@@
To directly compare SN and AGN feedback within the same framework, in Figure 6 we plot their respective contributions to the total gas expelled as a function of halo mass and redshift  [^3] In our model only SN and AGN can eject gas.. SN feedback is represented by solid lines, and AGN feedback by dashed lines, with different colors indicating redshift. The left and right panels correspond to the AGNeject1 and AGNeject2 models, respectively. The two models display clearly distinct behaviors: AGNeject1 shows a weak dependence on redshift, whereas AGNeject2 exhibits a strong redshift dependence, as expected from its delayed onset of AGN activity.
Notably, in both models, SN feedback dominates the total gas expulsion except within the characteristic mass ranges where AGN feedback becomes particularly efficient. Specifically, AGN feedback dominates in halos with $12.5<\log M_{\rm{halo}}<13.5$ for AGNeject1, and $12.2<\log M_{\rm{halo}}<13.3$ for AGNeject2, both at $z=0$. Thus, SN feedback is the primary driver of gas ejection in low-mass and Milky Way-like halos, while AGN feedback takes over in the subsequent mass decade, with both mechanisms contributing roughly equally on large group and cluster scales. It is important to note that in our SAM only SN and AGN feedback are capable of ejecting gas. On cluster scales, their contributions are comparable, although the absolute amount of expelled gas can be relatively small. We also remind the reader that the plot accounts for all progenitors in the merger trees, not just the main branch.
Our findings are qualitatively consistent with those of M. Ayromlou et al. (2023) (but see also the analysis in R. J. Wright et al. 2024), who analyzed three sets of cosmological hydrodynamic simulations—EAGLE (J. Schaye et al., 2015), IllustrisTNG (A. Pillepich et al., 2018), and SIMBA (R. Davé et al., 2019). These authors showed that feedback processes play a crucial role in redistributing gas within and beyond halo boundaries, reducing the baryon fraction inside halos and transferring hot gas outside the virial radius.
@@
To conclude this section, we now turn to a detailed comparison between the baryon fractions predicted by our models and those reported by M. Ayromlou et al. (2023) in their analysis.